2 blocks of masses 10 kg and 20 kg moving at speeds of 10m/sec and 20m/sec resp. in oppsite directions collide.If the collision is completly inelastic,find the thermal energy developed in the process.

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Allina,

Ans;- The final velocity of the combination is v (let) towards the initial direction of the 20 kg body Hence the Eq of the conservation of momentum is

(20*20)-(10*10)= 30v

Hence v=10m/s

From the conservation of the total energy we get,

1/2(20*20²+10*10²)=H+1/2(30*10²)

Solving we get H=2000 J

                       =2 kJ

Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

All the best Alina!!!

Regards,

Askiitians Experts SOUMYAJIT IIT_KHARAGPUR